# Do assumptions about the central density of subhaloes affect dark matter annihilation and lensing calculations?

### v2

@@ Abstract
A growing body of evidence suggests that the central density of cuspy dark matter subhaloes is conserved in minor mergers. In contrast to this, empirical models of subhalo evolution, calibrated using limited-resolution simulations, often assume a drop in the central density. These empirical models are an important component in galaxy-galaxy lensing studies and dark matter annihilation calculations. We explore the dependence of these calculations on the assumed subhalo properties. We find that annihilation calculations are very sensitive to the assumed inner density profile, and different models can produce more than an order of magnitude difference in the $J$-factor and boost factor in individual haloes, and a factor of $\sim 5$ in the total annihilation rate expected in the halo Milky Way. Since the innermost parts of haloes will always be difficult to resolve in simulations, we conclude that it is important to develop a theoretical understanding of subhalo evolution to be able to make accurate predictions of the dark matter annihilation signal. On the other hand, while the shear and convergence profiles used in galaxy-galaxy lensing are sensitive to the initial profile assumed (e.g., NFW versus Einasto), they are otherwise well-approximated by a simple stripping model in which the original profile is sharply truncated at a tidal radius.
@@ 1 Introduction
Dark matter haloes grow through repeated, hierarchical mergers, and simulations of this process predict that the central regions of merging haloes can survive as self-bound substructures within the final system. These ‘subhaloes’ should host most of the visible galaxies in the low-redshift Universe, and may be one of the best environments in which to explore and constrain the particle nature and non-gravitational properties of dark matter [see e.g. 7, for a review].
@@
In general, observational tests of this predicted dark matter distribution include galaxy dynamics, gravitational lensing, and ‘indirect detection’ (i.e., searches for radiation or products from dark matter annihilation). All of these tests are most sensitive where dark matter densities are highest, at small radii within haloes and subhaloes. The annihilation signal in particular will depend sensitively on subhalo density profiles, concentration, mass loss, and disruption [as summarized in, e.g. 68, 1]. Since this signal is proportional to the local density squared, one must understand the details of the density distribution within haloes and subhaloes, down to the smallest scales on which CDM can cluster, to place reliable constraints on dark matter particle properties.
@@
The visible structure of galaxies can be used to trace the central density distribution in haloes [e.g. 81, and references therein], but only out to a few percent of the virial radius [53, 75]. Typically, measurements of the inner core or cusp of dwarf-galaxy haloes require tens of thousands of stars and are sensitive to dynamical assumptions [9]. Gravitational lensing studies can probe the total mass distributions around galaxies, groups, and clusters more directly. However, while gravitational lensing has the advantage of being insensitive to the dynamical state of the system, the effect is normally weak enough that good models of the lens potential are required. As first shown by [61] and since measured in several studies [e.g. 56, 65, 74, 30, 54, 85], the tidal truncation of galaxy haloes within clusters relative to those of field galaxies is detectable in lensing and needs to be included in models of the total mass distribution to obtain accurate halo mass estimates [3].
@@
As subhaloes orbit within a larger system post-merger, they will lose mass through tidal stripping. This process generally works from the outside in [e.g. 39, 23]. However, the exact effect of tidal stripping in the innermost part of the subhalo remained unclear from previous work, given the resolution limits of simulations. In principle, repeated mass loss can also disrupt substructure completely, as is often seen in cosmological simulations of halo formation. Recent work suggests that much of this disruption is artificial, and due to insufficient resolution [83, 82, 5], and that the central density of cuspy subhaloes should be preserved to arbitrarily small masses [50, 33, 27].
@@
Fundamentally, subhalo structure depends on the initial properties of the halo at infall and the subsequent effect of tidal evolution. Studies of the earliest forming haloes—which are expected to exist as subhaloes at low redshifts—suggest that these objects are cuspier than systems that form later, and have an inner density profile of the form $r^{-1.5}$ [e.g. 45, 2, 67, 15, 19] [^2] This is steeper than the commonly-used NFW profile, which has an inner density profile of $r^{-1}$, as discussed in Section 2.1. Given the strong dependence of the annihilation signal on the dark matter density on the smallest scales and in the densest systems, this has enormous implications for current annihilation constraints [16]. For instance, it has been suggested that there should be $\sim 10^{16}$ “prompt cusps" in the Milky Way [18], which would contribute 20-80 per cent of a putative annihilation $\gamma$-ray background [79].
@@
The current study aims to consider how modelling assumptions for subhalo density profiles affect the dark matter annihilation and lensing signals in individual subhaloes. The structure of this paper is as follows: first, in Section 2, we summarize the subhalo models considered in this work. In Section 3, we show how the concentration evolves in these different models. In Sections 4 and 5, we explore the implications of a conserved central density to the dark matter annihilation rate and lensing signal, respectively. Finally, we discuss our results in Section 6 and conclude in Section 7.
@@ 2.1 Isolated dark matter haloes
where $\rho_{-2}$ and $r_{-2}$ correspond to where the logarithmic slope is ${\rm d}\log\rho/{\rm d}\log r=-2$. This fit provides a better description of the UDP down to the resolution limit of simulations [64, 35, 52]; the detailed form of the profile at very small radii remains uncertain.
The “shape” parameter, $\alpha$, controls the inner slope of the density profile, with small $\alpha$ values corresponding to “cuspier" centres. The mean value of $\alpha$ increases with peak height within the cosmological density field [35, 52]; for low peaks (i.e. low-mass and/or low redshift haloes) with values of $\alpha\sim$ 0.2–0.25, the Einasto profile is very similar to the NFW profile, while for rare, recently-formed and/or high-mass objects, it is smoother with a shallower inner slope.
Aside from their density profiles, haloes are often described by their mass and concentration. The concentration parameter $c$ is usually defined as the ratio of the virial radius to the scale radius, or more generally, the radius at which the logarithmic density profile has a slope of $-2$, i.e., $c=r_{\rm vir}/r_{-2}$ [63]. Concentration is broadly correlated with formation epoch, early-forming haloes being more concentrated, though the details of the relationship are complicated [e.g. 87, 92, 91, 89, 58, 11, 86].
@@ 2.2 Tidally-stripped dark matter subhaloes
In H03, the bound mass fraction $f_{\rm b}$ is defined as the mass of the bound satellite compared to the mass of an untruncated NFW profile within radius $r_{\rm vir}=10$.
The disadvantage of empirical methods like H03 is that these models are only valid across the range of models considered in the simulations, and therefore are generally limited to specific density profiles and orbital parameters. Additionally, these models may include numerical artefacts. For example, H03 predicts an artificial reduction in the central density due to the approximation they used to set up the initial conditions in their simulations [49]. More recent empirical models [e.g. 69, 36] predict similar trends but with less reduction in the central density. In general, empirical models accurately reproduce a suite of isolated simulation results by construction, but cannot be reliably extrapolated beyond the suite of simulations. Since the very centre of haloes will never be satisfactorily resolved in $N$-body simulations due to two-body interactions, these parameterizations are not able to predict the evolution at small radii/small mass fractions [27].
An alternative approach is to model the evolution of subhaloes using physical principles. This approach is generally less accurate than calibrated empirical models, with the possible exception being the Energy Truncation model developed in [25, 26] [see also 88]. The Energy Truncation model is based on the observation that particles are primarily stripped as a function of their energy [e.g. 10, 25, 78]. In practice, this energy truncation is performed by lowering and shifting the distribution function, $f_{0}$, of the initial profile, according to
@@
where $\mathcal{E}=\Psi(r)-v^{2}/2$ is the relative “binding" energy and $\Psi(r)=-\phi(r)$ is the relative potential energy. The parameter $\mathcal{E}_{T}$ is termed the truncation energy and sets the mass and tidal radius of the truncated system. Then, the potential of this system can be found by solving Poisson’s equation and Eddington’s inversion of the density profile [31]. This procedure is analogous to the derivation of the King model [51] and has been described in detail in [25, 26, 27]. The free parameter $\mathcal{E}_{T}$ can be determined from the orbital parameters of the merger [26], and thus, the Energy Truncation model has no free parameters (i.e., it does not need to be tuned to simulations).
Figure 1 demonstrates how well the Energy Truncation model predicts the subhalo density profile evolution compared to empirical models calibrated to match simulations [39, 69, 36]. We have used two simulations; Example 1 is demonstrative of a case in which the Energy Truncation Model works quite well, and Example 2 is a case where the Energy Truncation Model does not work as well; these correspond to the Slow and Fast Mass loss simulations in [27]. In general, the Energy Truncation model predicts a higher central density than the other models, and the [39] model (H03) predicts the lowest central density.
Figure 1: Comparison of models for tidally stripped NFW haloes. The simulation results are from an idealized simulation of an NFW subhalo inside a fixed background potential after 5 orbits. Examples 1 and 2 correspond to the Slow and Fast Mass loss simulations in [27]. The Energy Truncation Model [26] has no free parameters, while the three parametric models [39, 69, 36] are calculated from the bound mass of the simulated subhalo. Overall, all of the models do a comparable job of predicting the subhalo density profile.
Overall, the Energy Truncation model agrees very well with isolated simulations and has comparable accuracy to parameterized models above within the range of radii that is properly resolved in simulations [26]. Further, the Energy Truncation model appears to be universally applicable to any tidally-stripped collisionless system [27]. Unlike most empirical parameterizations, the Energy Truncation model predicts the central density is preserved, which may have important consequences for applications such as dark matter annihilation and lensing predictions.
@@ 2.3 This work
Overall there is some fundamental uncertainty in the original density profile of isolated haloes (e.g. NFW and the range of Einasto profiles) and the way they are modified by tidal stripping. It is impossible to completely resolve these uncertainties by direct numerical simulation: given the limited dynamic range of simulations, we will never be able to resolve the very centre of haloes. As a result, the central density of subhaloes may be higher than previously expected. In the remaining sections of the paper, we examine how this influences the dark matter annihilation and galaxy lensing signals.
We will consider three initial profiles
@@
    a more “cuspy" Einasto profile (EinLow; $\alpha=0.15$), and
@@
    a more “cored" Einasto profile (EinHigh; $\alpha=0.3$).
These Einasto parameters were chosen since they span the range of parameters found in simulations [e.g. 35, 52]. The lower value is typical of low peaks, i.e. late-forming or low-mass objects, while the high value is typical of early-forming or high-mass objects. For comparison, a commonly adopted fiducial Einasto parameter is 0.17 [e.g. 55]. The NFW profile should be intermediate, and most similar to an Einasto profile with $\alpha\sim 0.2$–0.25. We will use units $G=1$, $r_{\rm unit}=r_{-2}$, and $M_{\rm unit}=M_{\rm NFW}(r<10\,r_{-2})$; i.e., of profiles will be normalized to have the same virial mass and a concentration of $c=10$. As shown in Figure 2, though these profiles have the same virial mass and concentration, at radii less than 10 per cent of scale radius, the central densities begin to differ considerably.
@@
We will primarily use the Energy Truncation method to describe our tidally stripped systems and compare these results to two commonly used approximations in the literature: (1) the H03 model and (2) a sharp truncation (ST) of an NFW model. The H03 model is known to underestimate the central density, while the ST model preserves the NFW profile precisely within the tidal radius. These two approximations of a tidally-stripped NFW profile will therefore serve as limiting cases. The parameterizations by [69] and [36] are expected to fall somewhere between the Energy Truncation model and the H03 model.
@@
    Differences in predictions caused by assumptions about the initial profile model; i.e., between NFWT, EinHighT, and EinLowT
@@
Table 1: Summary of the subhalo models used in this work. For all models, subhaloes are assumed to have an infall mass of $M_{\rm unit}=M_{\rm NFW}(r<10\,r_{-2})$ and a concentration of $c=10$.
@@
EinHighT EinHigh Energy Truncation
EinLowT EinLow Energy Truncation
@@
Figure 3: Tidally stripped subhalo density profiles. The models are summarized in Table 1. Each panel is tripped to a different bound mass. As in Figure 2, the grey box indicates radii that are beyond what is typically resolved in isolated simulations. The Energy Truncation model predictions (NFWT, EinHighT, and EinLowT) are similar, except at very small radii. The two other stripping models differ significantly from the Energy Truncation model. The H03 model underestimates the density of the subhaloes at small mass fractions, while the ST model only agrees with NFWT at small radii.

### v3

@@ Abstract
Subhalo models play a critical role in dark matter annihilation predictions and galaxy-galaxy lensing studies; however the internal structure of subhaloes remains highly uncertain. In particular, a growing body of evidence suggests that the central density of cuspy dark matter subhaloes is conserved in minor mergers, whereas empirical models of subhalo evolution—calibrated using limited-resolution simulations—often assume a drop in the central density. To assess the impact of these assumptions, we systematically explore how a wide range of initial mass profiles and tidal evolution prescriptions influence annihilation and lensing calculations, including the physically motivated Energy Truncation model, which explicitly preserves the central density of subhaloes. We find that annihilation calculations are very sensitive to the assumed inner density profile, and different models can produce more than an order of magnitude difference in the annihilation rate of individual subhaloes, and a factor of $\sim 5$ in the total annihilation rate expected in the Milky Way. Since the innermost parts of haloes will always be difficult to resolve in simulations, we conclude that developing a theoretical understanding of subhalo evolution is crucial to be able to make accurate predictions of the dark matter annihilation signal. On the other hand, while the shear and convergence profiles used in galaxy-galaxy lensing are sensitive to the initial profile assumed (e.g., NFW versus Einasto), they are otherwise well-approximated by a simple stripping model in which the original profile is sharply truncated at a tidal radius.
@@ 1 Introduction
Dark matter haloes grow through repeated, hierarchical mergers, and simulations of this process predict that the central regions of merging haloes can survive as self-bound substructures within the final system. These subhaloes should host most of the visible galaxies in the low-redshift Universe, and may be one of the best environments in which to explore and constrain the particle nature and non-gravitational properties of dark matter [see e.g. 1, for a review].
@@
In general, observational tests of this predicted dark matter distribution include galaxy dynamics, gravitational lensing, and ‘indirect detection’ (i.e., searches for radiation or products from dark matter annihilation). All of these tests are most sensitive where dark matter densities are highest, at small radii within haloes and subhaloes. The annihilation signal in particular will depend sensitively on subhalo density profiles, concentration, mass loss, and disruption [as summarized in, e.g. 2, 3]. Since this signal is proportional to the local density squared, one must understand the details of the density distribution within haloes and subhaloes, down to the smallest scales on which CDM can cluster, to place reliable constraints on dark matter particle properties.
@@
The visible structure of galaxies can be used to trace the central density distribution in haloes [e.g. 4, and references therein], but only out to a few per cent of the virial radius [5, 6]. Typically, measurements of the inner core or cusp of dwarf-galaxy haloes require tens of thousands of stars and are sensitive to dynamical assumptions [7]. Gravitational lensing studies can probe the total mass distributions around galaxies, groups, and clusters more directly. However, while gravitational lensing has the advantage of being insensitive to the dynamical state of the system, the effect is normally weak enough that good models of the lens potential are required. As first shown by [8] and since measured in several studies [e.g. 9, 10, 11, 12, 13, 14], the tidal truncation of galaxy haloes within clusters relative to those of field galaxies is detectable in lensing and needs to be included in models of the total mass distribution to obtain accurate halo mass estimates [15].
@@
As subhaloes orbit within a larger system post-merger, they will lose mass through tidal stripping. This process generally works from the outside in [e.g. 16, 17]. However, the exact effect of tidal stripping in the innermost part of the subhalo has remained unclear from previous work, given the resolution limits of simulations. In principle, repeated mass loss can also disrupt substructure completely, as is often seen in cosmological simulations of halo formation. Recent work suggests that much of this disruption is artificial, and due to insufficient resolution [18, 19, 20, 21], and that the central density of cuspy subhaloes should be preserved to arbitrarily small masses [22, 23, 24, 25].
@@
Fundamentally, subhalo structure depends on the initial properties of the halo at infall and the subsequent effect of tidal evolution. Studies of the earliest forming haloes—which are expected to exist as subhaloes at low redshifts—suggest that these objects are cuspier than systems that form later, and have an inner density profile of the form $r^{-1.5}$ [e.g. 26, 27, 28, 29, 30] [^2] This is steeper than the commonly-used NFW profile, which has an inner density profile of $r^{-1}$, as discussed in Section 2.1. Given the strong dependence of the annihilation signal on the dark matter density on the smallest scales and in the densest systems, this has enormous implications for current annihilation constraints [31]. For instance, it has been suggested that there should be $\sim 10^{16}$ “prompt cusps" in the Milky Way [32], which would contribute 20-80 per cent of a putative annihilation $\gamma$-ray background [33].
@@
The current study investigates how different modelling assumptions for subhalo density profiles impact dark matter annihilation and lensing signals in individual subhaloes. In particular, we focus on the consequences of assuming that the central density of cuspy subhaloes is preserved during tidal evolution. The structure of this paper is as follows: first, in Section 2, we summarize the subhalo models considered in this work. In Section 3, we show how the concentration evolves in these different models. In Sections 4 and 5, we explore the implications of a conserved central density to the dark matter annihilation rate and lensing signal, respectively. Finally, we discuss our results in Section 6 and conclude in Section 7.
@@ 2.1 Isolated dark matter haloes
where $\rho_{-2}$ and $r_{-2}$ correspond to where the logarithmic slope is ${\rm d}\log\rho/{\rm d}\log r=-2$. Though this fit provides a better description of the UDP down to the resolution limit of simulations [37, 38, 39], the detailed form of the profile at very small radii remains uncertain.
The “shape” parameter, $\alpha$, controls the inner slope of the density profile, with small $\alpha$ values corresponding to “cuspier" centres. The mean value of $\alpha$ increases with peak height within the cosmological density field [38, 39]; for low peaks (i.e. low-mass and/or low redshift haloes) with values of $\alpha\sim$ 0.2–0.25, the Einasto profile is very similar to the NFW profile, while for rare, recently-formed and/or high-mass objects, it is smoother with a shallower inner slope. Aside from their density profiles, haloes are often described by their mass and concentration. We discuss the definition of the evolution of the concentration parameter in Section 3.
@@ 2.2 Tidally-stripped dark matter subhaloes
In H03, the bound mass fraction $f_{\rm b}$ is defined as the mass of the bound satellite compared to the mass of an untruncated NFW profile within radius $r_{\rm vir}=10\,r_{s}$.
The disadvantage of empirical methods like H03 is that these models are only valid across the range of cases considered in the simulations and are, therefore, generally limited to specific density profiles and orbital parameters. Additionally, these models may include numerical artefacts. For example, H03 predicts a reduction in the central density, which is a consequence of an artefact of the approximation they used to set up the initial conditions in their simulations [40]. More recent empirical models [e.g. 41, 42] predict similar trends but with less reduction in the central density. In general, empirical models accurately reproduce a suite of isolated simulation results by construction, but cannot be reliably extrapolated beyond the suite of simulations. Since the very centre of haloes will never be satisfactorily resolved in $N$-body simulations due to two-body interactions, these parameterizations are not able to predict the evolution at small radii/small mass fractions [25].
An alternative approach is to model the evolution of subhaloes using physical principles. This approach is generally less accurate than calibrated empirical models, with the possible exception being the Energy Truncation model developed in [43, 44]; see also [45]. The Energy Truncation model is based on the finding that particles are primarily stripped as a function of their energy [e.g. 46, 43, 47]; particles that are less gravitationally bound get stripped first, regardless of their radius.
In practice, this energy truncation is performed by lowering and shifting the distribution function, $f_{0}$, of the initial profile, according to
@@
where $\mathcal{E}=\Psi(r)-v^{2}/2$ is the relative “binding" energy and $\Psi(r)=-\phi(r)$ is the relative potential energy. The parameter $\mathcal{E}_{T}$ is termed the truncation energy and sets the mass and tidal radius of the truncated system. Then, the potential of this system can be found by solving Poisson’s equation and Eddington’s inversion of the density profile [48]. This procedure is analogous to the derivation of the King model [49] and has been described in detail in [43, 44, 25]. The free parameter $\mathcal{E}_{T}$ can equivalently be expressed as a bound mass or tidal radius. In this work we primarily express subhaloes by their bound mass, but, in principle, $\mathcal{E}_{T}$ can be determined from the orbital parameters of the merger [44], and thus, the Energy Truncation model has no free parameters (i.e., it does not need to be tuned to simulations).
Figure 1 demonstrates how well the Energy Truncation model predicts the subhalo density profile evolution compared to empirical models calibrated to match simulations [16, 41, 42]. While empirical models are calibrated to match simulation results, the impact of numerical artefacts on the central density profile remains uncertain. The Energy Truncation model predicts slightly higher central densities than the empirical models but agrees well with isolated simulations and provides comparable accuracy to parameterized models within the radii that are well resolved in simulations [44]. Additionally, this model has been shown to reproduce not only the density profile but also the phase-space structure of tidally stripped systems undergoing gradual mass loss, particularly when evaluated near apocentre [43]. However, we do not expect the Energy Truncation Model to accurately describe all scenarios—for instance, it may not perform well in systems experiencing rapid or impulsive stripping (e.g., disk shocking), where energy redistribution can be more complex.
Figure 1: Comparison of models for tidally stripped NFW haloes. The simulation results are from an idealized simulation of an NFW subhalo inside a fixed background potential after 5 orbits. We have used two simulations; Example 1 is demonstrative of a case in which the Energy Truncation Model works quite well, and Example 2 is a case where the Energy Truncation Model performs less accurately; these correspond to the Slow and Fast Mass loss simulations in [25]. The Energy Truncation Model [44] has no free parameters, while the three parametric models [16, 41, 42] are calculated from the bound mass of the simulated subhalo. Overall, all of the models do a comparable job of predicting the subhalo density profile.
Since the Energy Truncation model predicts that the central density is preserved—unlike most empirical parameterizations—we adopt this model as a physically motivated upper bound on central densities. Additionally, because the Energy Truncation model is applicable to any tidally stripped collisionless system [25], it offers a flexible approach for studying different profile models without requiring calibration to simulations.
@@ 2.3 This work
Overall there is some fundamental uncertainty in the original density profile of isolated haloes (e.g. NFW and the range of Einasto profiles) and the way they are modified by tidal stripping. It is impossible to completely resolve these uncertainties by direct numerical simulation: given the limited dynamic range of simulations, we will never be able to resolve the very centre of low-redshift haloes in any near future. As a result, the central density of subhaloes may be higher than previously expected. In the remaining sections of the paper, we examine how conserved central density and varying initial profiles influence the dark matter annihilation and galaxy lensing signals.
We will consider three initial profile models
@@
    a more “cuspy" Einasto profile (EinCusp [^3] In [25], EinCusp and EinCore were called EinLow and EinHigh, respectively.; $\alpha=0.15$), and
@@
    a more “cored" Einasto profile (EinCore; $\alpha=0.3$).
These Einasto parameters were chosen since they span the range of parameters found in simulations [e.g. 38, 39]. The lower value is typical of low peaks, i.e. late-forming or low-mass objects, while the high value is typical of early-forming or high-mass objects. For comparison, a commonly adopted fiducial Einasto parameter is 0.17 [e.g. 50]. The NFW profile should be intermediate between the two, and most similar to an Einasto profile with $\alpha\sim 0.2$–0.25. We will use units $G=1$, $r_{\rm unit}=r_{-2}$, and $M_{\rm unit}=M_{\rm NFW}(r<10\,r_{-2})$; that is, the profiles will be normalized to have the same virial mass and a concentration of $c=10$. The density unit is then $\rho_{\rm unit}=M_{\rm unit}r_{\rm unit}^{-3}$. As shown in Figure 2, although these three profiles have the same virial mass and concentration, at radii less than 10 per cent of scale radius, the central densities begin to differ considerably.
@@
We will primarily use the Energy Truncation model to describe our tidally stripped systems and compare these results to two commonly used approximations in the literature: (1) the H03 model and (2) a sharp truncation (ST) of an NFW model. The H03 model is known to underestimate the central density, while the ST model preserves the NFW profile precisely within the tidal radius; therefore, the Energy Truncation model can be used as an intermediate, physically-grounded alternative to both the H03 (which suffers from numerical limitations) and the sharp truncation model (which does not account for any internal structural change). The parameterizations of [41] and [42] are expected to fall somewhere between the Energy Truncation model and the H03 model. Given this, we adopt the Energy Truncation model as a conservative upper bound on subhalo density for cuspy systems, representing a scenario in which the central density is fully preserved.
@@
    Differences in predictions caused by assumptions about the initial profile model under the hypothesis of a conserved central density; i.e., between NFWT, EinCoreT, and EinCuspT
@@
Table 1: Summary of the subhalo models used in this work. For all models, subhaloes are assumed to have an initial infall mass of $M_{\rm unit}=M_{\rm NFW}(r<10\,r_{-2})$ and a concentration of $c=10$. This leaves one free parameter; the bound mass, $M$, of the satellite.
@@
EinCoreT EinCore Energy Truncation
EinCuspT EinCusp Energy Truncation
@@
Figure 3: Tidally stripped subhalo density profiles. The models are summarized in Table 1. Each panel is stripped to a different bound mass. As in Figure 2, the grey box indicates radii that are beyond what is typically resolved in isolated simulations. The Energy Truncation model predictions (NFWT, EinCoreT, and EinCuspT) are similar, except at very small radii. The two other stripping models differ significantly from the Energy Truncation model. The H03 model is known to underestimate the density of the subhaloes at small mass fractions, while the ST model only agrees with NFWT at small radii.